# bibliography.bib

### v1

@@ A.1 Linear IV
    Confounding changes the query: a confounded SCM $\mathcal{S}_{1}$ and an unconfounded $\mathcal{S}_{2}$ both compatible with $\mathcal{C}$ cannot coincide in their observational distributions, i.e., $\mathbb{P}_{\mathrm{obs}}^{S_{1}}\neq\mathbb{P}_{\mathrm{obs}}^{S_{2}}$.
Point 1) is a classical positivity assumption as common in the causal inference literature. Point 2) ensures that an unconfounded DGP is a possibility when generating the observational data. Finally, point 3) ensures that causal query $Q$ is affected by the unobserved confounder $U$ between $A$ and $Y$. If 3) would not hold, we find an unconfounded $\mathcal{S}_{1}$ and and unconfounded $\mathcal{S}_{2}$ with $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{1}}=\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{2}}$, which also implies $Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{1}})=\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{1}})=\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{2}})=Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{2}})$ by identifiability of the causal query. Hence, this assumption is realistic for various causal queries, such as CAPO or CATE.
Proof.
Let $\Pi$ be a well-specified $\mathcal{C}$-SCM-prior with induced C-DAG $G_{c}$ that has an unobserved confounder between $A$ and $Y$. We have to show that in $G_{c}$ either $A$ or $Y$ has a noise parent.
By Assumption A.1 (2), there exists SCM $\mathcal{S^{\ast}}$ compatible with $\mathcal{C}$, in which $A$ and $Y$ are unconfounded. Let $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}^{\ast}}$ be sampled from the unconfounded observational distribution. Because $\Pi$ is well specified for $\mathcal{C}$, it holds that
$Q\!\Bigl{(}\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}\Bigr{)}\;\longrightarrow\;Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}^{\ast}}),\quad n\to\infty.$ (63)
However, this must imply that for all $n\in\mathbb{N}$ there exists an SCM $\mathcal{S}$ with posterior $\Pi(\mathcal{S}\mid\mathcal{D}_{n})>0$, and hence $\Pi(\mathcal{S})>0$ and $\Pi(\mathcal{D}_{n}\mid\mathcal{S})>0$. Thus in the limit for $n\to\infty$ this implies $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}=\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}^{\ast}}$. By assumption  A.1 (3) this implies that $\mathcal{S}$ is unconfounded. Hence, because of Assumption A.1 (1), either $A$ or $Y$ must have a noise parent in $\mathcal{S}$. Because $\Pi(\mathcal{S}>0)$, either $A$ or $Y$ must have also a noise parent in $G_{c}$.
Appendix B Implementation details
We implement our CausalFM using PyTorch. Our transformer-based model is designed for conditional average treatment effect (CATE) estimation. Our model implementation builds upon the TabPFN architecture [Hollmann.2023] https://github.com/PriorLabs/TabPFN/tree/main. We use separate encoders for each modality. Each observation in the dataset is tokenized during the embedding process. The embedded tokens are then processed through a basic transformer where attention is applied between the observations. In the final layer, tokens corresponding to the inference observations are passed through a classification head. We frame CATE estimation as a three-class classification task by discretizing the individual treatment effect into three categories: negative effect, no effect, and positive effect. The model uses a task-specific decoder to predict CATE. During training, we use cross-entropy loss computed on model output and target classes. When training from scratch, the model is optimized using the Adam optimizer with a learning rate of $31e-$ 3 and weight decay of $51e-$ 5 .. The training batch size is set to $16$ and the sequence length is set to $1024$. The total training epoch is $100$. Early stopping is triggered after $10$ epochs of no improvement in validation loss. We generate $10$ k datasets. Empirically, training our causalFM model takes about $10$ hours on an NVIDIA A100 GPU.
B.1 Implementation details of baselines
We follow the implementation from https://github.com/AliciaCurth/CATENets/tree/main for most of the CATE estimators, including S-learner [kunzel2019metalearners], T-learner [kunzel2019metalearners], TARNet [shalit2017estimating], X-leaner [kunzel2019metalearners], DR-learner [kennedy2023towards], RA-learner [curth2021nonparametric]. For each dataset and method, we evaluated 5 repetitions, each with a different random seed. All methods used the same train-test split.

### v2

@@ E.6 Generation pipeline
    For outcomes, sample $U_{Y}$ once per unit and use it to compute $Y(0)$ and $Y(1)$ via (62).
E.7 Synthetic datasets size
We sample $10000$ synthetic training datasets from data prior with different data generation mechanism. Each training datasets contain $1024$ data samples. The feature dimensions are also different across the datasets, ranging from $10$ to $100$. The features are mixed data type with continuous, binary and categorical.
Appendix F Synthetic data generation for the instrumental variables (IV) setting
We aim at estimating CATEs from observational data under unobserved confounding using IVs. In contrast to the standard CATE setting, where strong unconfoundedness holds, our IV datasets intentionally violate unconfoundedness by introducing an unobserved confounder $U$ that affects both treatment $A$ and outcome $Y$. Identification is instead driven by an instrument $Z$ that (i) is relevant for $A$, (ii) has no direct path to $Y$ beyond $A$ (exclusion), and (iii) is conditionally independent of $U$ given $X$.
Key differences vs. standard CATE.
(i) Ignorability is broken: $A\not\!\perp\!\!\!\perp\{Y(0),Y(1)\}\mid X$ due to $U\to A$ and $U\to Y$. (ii) We introduce an instrument $Z$ with $Z\perp U\mid X$, $Z\not\!\perp\!\!\!\perp A\mid X$, and no $Z\to Y$ edge (exclusion). (iii) Outcomes are generated via an additive structural form $Y=f(X,A)+g(X,U)+\varepsilon_{Y}$ with $f$ and $g$ deterministic neural networks; the same $\varepsilon_{Y}$ is reused across $Y(0)$ and $Y(1)$ for a unit to ensure counterfactual consistency.
F.1 Covariates and latent confounders via a DAG-structured SCM
We reuse the DAG-SCM from the standard setting to produce a wide set of base variables $W$, then split it into observed covariates $X$ and unobserved confounders $U$. Thus we have different strength of the unobserved confounders from weak to sufficiently strong.
Graph and node equations.
Sample number of layers $L_{X}$ and hidden size $H_{X}$, build a layered DAG (fully connect layer $\ell$ to $\ell{+}1$), and drop a fraction $p_{\text{drop}}^{X}$ of inter-layer edges to sparsify. For each node $j$, sample weights $\{w_{jk}\}_{k\in\mathrm{pa}(j)}$, bias $b_{j}$, and a node-specific exogenous noise $\varepsilon_{j}\sim\mathcal{D}_{j}$ (type and scale drawn once per node). Evaluate in topological order
$s_{j}\;=\;\sum_{k\in\mathrm{pa}(j)}w_{jk}\,v_{k}\;+\;b_{j}\;+\;\varepsilon_{j},\qquad v_{j}\;=\;\tanh(s_{j}).$ (65)
[…]
F.2 Instrument variable
We generate $Z$ from $X$ only, ensuring $Z\perp U\mid X$ by construction and precluding any direct $U\to Z$ path. Let $\phi_{Z}$ be a feed-forward network with input $X$ and no layer-wise exogenous noise; the network parameters are sampled once per dataset and then fixed. For a unit with covariates $X$,
$s_{Z}\;=\;\phi_{Z}(X),\qquad Z\;=\;\begin{cases}\mathrm{Bernoulli}\!\big(\sigma(s_{Z})\big),&\text{binary instrument},\\[2.0pt] s_{Z},&\text{continuous instrument}.\end{cases}$ (66)
We randomly choose between the binary and continuous variants when creating datasets. Relevance is induced via the $Z\to A$ path in the treatment mechanism below.
Note that the instrument variable $Z$ has a direct influence on the treatment $A$, but does not have a direct effect on the outcome $Y$.
F.3 Treatment variable
Given $(X,Z,U)$, treatment is generated via a deterministic network $\phi_{A}$ followed by a Bernoulli draw. There is no layer-wise noise inside $\phi_{A}$; the only randomness is the terminal Bernoulli. For a unit,
$s_{A}\;=\;\phi_{A}\big([X;Z;U]\big),\qquad p\;=\;\sigma(s_{A}),\qquad A\sim\mathrm{Bernoulli}(p).$ (67)
This introduces $U\to A$ and hence breaks ignorability, while maintaining $Z\perp U\mid X$ and $Z\to A$ relevance.
F.4 Outcome variables
The instrument variable $Z$ has no direct effect on the outcomes. Outcomes are generated additively from a treatment channel $f$ and a confounding channel $g$, both deterministic MLPs with inputs $[X;A]$ and $[X;U]$, respectively. Let $\varepsilon_{Y}\sim\mathcal{N}(0,\sigma_{Y}^{2})$ be an i.i.d. scalar noise drawn once per unit,
$Y(a)\;=\;f\big(X,a\big)\;+\;g(X,U)\;+\;\varepsilon_{Y},\qquad a\in\{0,1\},$ (68)
Y=AY(1)+(1A)Y(0).Y\;=\;A\,Y(1)+(1-A)\,Y(0). (69)
By construction there is no $Z\to Y$ edge (exclusion), since $Z$ influences $Y$ only through $A$.
F.5 Independence and identification (IV)
All exogenous noises are sampled independently across units and mechanisms. The IV conditions hold by construction,
(Independence) $Z\perp U\mid X,$ (70)
(Exclusion) $Y(a)\;\text{depends on $X,a,U$ and $\varepsilon_{Y}$ only (no $Z$)},$ (71)
(Relevance) $Z\not\!\perp\!\!\!\perp A\mid X.$ (72)
F.6 Hyperparameters and priors
We use simple priors mirroring our implementation:
  • DAG-SCM for $(X,U)$: $L_{X}\sim\mathrm{Unif}\{2,3,4,5\}$,  $H_{X}\sim\mathrm{Unif}\{10,\dots,50\}$,  edge-drop $p_{\text{drop}}^{X}{=}0.4$; node noises $\varepsilon_{j}$ draw a type in {Normal, Uniform, Laplace, Logistic} with random scale.
  • Instrument net $\phi_{Z}$: depth $L_{Z}\!\geq\!3$,  width $H_{Z}\sim\mathrm{Unif}\{8,\dots,30\}$; output is either Bernoulli with $\sigma(s_{Z})$ (binary $Z$) or real-valued $s_{Z}$ (continuous $Z$); no layer-wise noise.
  • Treatment net $\phi_{A}$: depth $L_{A}\!\geq\!3$,  width $H_{A}\sim\mathrm{Unif}\{8,\dots,30\}$; no layer-wise noise; $A\sim\mathrm{Bernoulli}(\sigma(s_{A}))$.
  • Outcome nets $f,g$: depths $L_{f},L_{g}\sim\mathrm{Unif}\{3,\dots,6\}$,  widths $H_{f},H_{g}\sim\mathrm{Unif}\{10,\dots,25\}$; $\varepsilon_{Y}\sim\mathcal{N}(0,\sigma_{Y}^{2})$ with $\sigma_{Y}=0.5$ by default.
  • Weights/biases: i.i.d. $w,b\sim\mathcal{N}(0,1)$ sampled once per dataset; $\tanh$ activations.
  • Strength: $d_{U}\sim\mathrm{Unif}\{2,\dots,5\}$.
F.7 Generation pipeline (IV)
For each dataset, we execute:
  1. 1.
    Sample the covariate DAG and parameters; for each unit, evaluate (65) to obtain a wide matrix then split it into $(U,X)$.
  2. 2.
    Given $X$, compute the instrument $Z$ via (66) (binary or continuous and mixed).
  3. 3.
    Given $(X,Z,U)$, compute the treatment propensity $p=\sigma(s_{A})$ via (67) and sample $A\sim\mathrm{Bernoulli}(p)$.
  4. 4.
    Draw a single $\varepsilon_{Y}$ per unit and compute $Y(0),Y(1)$ using (68); set the factual outcome by (69).
This yields datasets matching the classical IV graph and enabling evaluation of IV estimators.
Appendix G Synthetic data generation for the Front-door–adjusted setting
G.1 Front-door adjustment datasets
We next construct datasets satisfying the front-door criterion. Besides covariates $X$, treatment $A$, and continuous outcomes $Y$, we introduce a mediator $M$. The design ensures that $A$ affects $Y$ only through $M$ (no direct $A\!\to\!Y$ path), $U$ (unobserved) confounds $A$ and $Y$ but does not affect $M$.
Covariates via a DAG-structured SCM.
Identical to the standard setting: we sample a layered DAG, draw node-wise weights/biases/noise, evaluate in topological order as in (56), and expose $d$ node values as observed features $X\in\mathbb{R}^{d}$. Independent exogenous noises $U_{X}{=}\{\varepsilon_{j}\}$ are re-sampled per unit.
Latent confounders.
From the same SCM evaluation we also retain $q$ additional node values as unobserved confounders $U\in\mathbb{R}^{q}$ (not revealed to learners). These induce confounding between $A$ and $Y$.
G.1.1 Treatment assignment with latent confounding
Given $(X,U)$, we sample a feed-forward network and generate treatment. Let $L_{A}\!\geq\!3$ and $H_{A}$ be the depth and width, respectively. With $h^{(0)}=[X,U]$,
$s_{A}^{(\ell)}\;=\;W_{A}^{(\ell)}h^{(\ell-1)}\;+\;b_{A}^{(\ell)},\quad h^{(\ell)}\;=\;\tanh\!\big(s_{A}^{(\ell)}\big),\;\;\ell=1,\dots,L_{A}-1,$ (73)
and scalar logit
$\tilde{s}_{A}\;=\;w_{A}^{\top}h^{(L_{A}-1)}\;+\;b_{A},\qquad p\;=\;\sigma(\tilde{s}_{A}),\qquad A\sim\mathrm{Bernoulli}(p).$ (74)
G.1.2 Mediator mechanism
The mediator is generated from $(X,A)$ only, thereby enforcing the front-door exclusion $U\nrightarrow M$. Let $L_{M}\!\geq\!3$, $H_{M}$ be depth and width, with input $g^{(0)}=[X,A]$,
$r^{(\ell)}\;=\;W_{M}^{(\ell)}g^{(\ell-1)}\;+\;b_{M}^{(\ell)}\;+\;\varepsilon_{M}^{(\ell)},\qquad g^{(\ell)}\;=\;\tanh\!\big(r^{(\ell)}\big),\;\;\ell=1,\dots,L_{M}-1,$ (75)
and scalar output
$M\;=\;w_{M}^{\top}g^{(L_{M}-1)}\;+\;b_{M}\;+\;\varepsilon_{M}^{(L_{M})}.$ (76)
We denote $U_{M}=\{\varepsilon_{M}^{(\ell)}\}_{\ell=1}^{L_{M}}$.
G.1.3 Outcome variable
Outcomes are constructed to satisfy $A\to M\to Y$ as the only causal path from $A$ to $Y$, while allowing $U\to Y$ and $X\to Y$. We decompose $Y$ into an $M$-path component and a confounding component:
[…]
$R(X,M)\;=\;v^{\top}h^{(L_{Y}-1)}\;+\;c\;+\;\xi^{(L_{Y})},$ (77)
Confounding path: $G(X,U)\;=\;\tilde{v}^{\top}\tilde{h}^{(L_{\!G}-1)}\;+\;\tilde{c}\;+\;\tilde{\xi}^{(L_{\!G})},\quad\tilde{h}^{(0)}=[X,U],\ \tilde{h}^{(\ell)}=\tanh(\cdot),$ (78)
and define the potential outcomes
[…]
The factual outcome is Y=AY(1)+(1A)Y(0)Y=A\,Y(1){+}(1{-}A)\,Y(0). By construction there is no direct $A\!\to\!Y$ edge; $A$ influences $Y$ solely via $M$.
G.1.4 Hyperparameters and priors
  • Covariate DAG: $L_{X}\sim\mathrm{Unif}\{3,4,5,6\}$, $H_{X}\sim\mathrm{Unif}\{15,\dots,40\}$, edge-drop $p_{\text{drop}}^{X}{=}0.5$; node noises drawn per-node from {Normal, Uniform, Laplace, Logistic} with random scale.
  • Treatment net (Eq. (73)–(74)): $L_{A}\sim\mathrm{Unif}\{3,4\}$, $H_{A}\sim\mathrm{Unif}\{8,\dots,20\}$.
  • Mediator net (Eq. (75)–(76)): $L_{M}\sim\mathrm{Unif}\{3,4\}$, $H_{M}\sim\mathrm{Unif}\{8,\dots,20\}$.
  • Outcome nets (Eq. (G.1.3)–(79)): $L_{Y},L_{\!G}\sim\mathrm{Unif}\{3,4,5\}$, widths $\sim\mathrm{Unif}\{10,\dots,25\}$; additive Gaussian $\epsilon_{Y}$ with task-specific scale.
  • Weights/biases: i.i.d. $\mathcal{N}(0,\sigma_{w}^{2})$; $\tanh$ nonlinearity.
G.1.5 Generation pipeline
For each dataset:
  1. 1.
    Sample the covariate DAG and evaluate to obtain $(X,U)$ (observed $X$, hidden $U$).
  2. 2.
    Compute $p(A{=}1\mid X,U)$ via (73)–(74) and sample $A$.
  3. 3.
    Evaluate the mediator $M$ from $(X,A)$ using (75)–(76).
  4. 4.
    Sample $U_{Y}$ once per unit and compute $Y(0)$ and $Y(1)$ via (79) by first obtaining $M(0)$ and $M(1)$ from the mediator net; set Y=AY(1)+(1A)Y(0)Y=A\,Y(1){+}(1{-}A)\,Y(0).
Appendix H Additional experiments
H.1 Additional results on the standard CATE estimation
We report the detailed standard CATE estimation on 10 synthetic datasets in Table H.1. We show our method gives the best estimation on most of the datasets.
Table 5: Standard CATE estimation on 10 synthetic datasets.
Method $\mathrm{D}_{1}$ $\mathrm{D}_{2}$ $\mathrm{D}_{3}$ $\mathrm{D}_{4}$ $\mathrm{D}_{5}$ $\mathrm{D}_{6}$ $\mathrm{D}_{7}$ $\mathrm{D}_{8}$ $\mathrm{D}_{9}$ $\mathrm{D}_{10}$
[…]
S-learner (Künzel et al., 2019) 0.725 0.583 0.752 0.829 0.614 0.892 0.858 0.421 0.680 0.985
T-learner (Künzel et al., 2019) 0.652 0.496 0.666 0.746 0.552 0.849 0.761 0.357 0.608 0.931
TARNet (Shalit et al., 2017b) 0.769 0.779 0.817 0.984 0.640 0.938 1.405 0.505 0.736 0.968
RA-learner (Curth & van der Schaar, 2021) 0.620 0.421 0.644 0.706 0.523 0.808 0.646 0.353 0.613 0.759
X-learner (Künzel et al., 2019) 0.574 0.400 0.614 0.634 0.381 0.713 0.686 0.302 0.549 0.779
DR-learner (Kennedy, 2023b) 0.783 0.533 0.767 0.947 0.867 0.882 0.791 0.4230 0.653 0.998
[…]
CausalPFN (Balazadeh et al., 2025) 0.493 0.489 0.585 0.743 0.413 0.615 0.950 0.288 0.453 0.544
DoPFN (Robertson et al., 2025) 0.417 0.313 0.228 0.679 0.591 0.475 0.497 0.551 0.610 0.827
CausalFM (ours) 0.454 0.487 0.515 0.677 0.204 0.618 0.950 0.278 0.442 0.532
Reported: PEHE (Lower = better, best in bold).
Table 6: Standard CATE estimation on ACIC2016 datasets. Reported: PEHE (mean $\pm$ std.)
Method PEHE
[…]
S-learner (Künzel et al., 2019) 1.191 $\pm$ 0.15
T-learner (Künzel et al., 2019) 1.143 $\pm$ 0.14
TARNet (Shalit et al., 2017b) 0.934 $\pm$ 0.15
RA-learner (Curth & van der Schaar, 2021) 0.762 $\pm$ 0.14
X-learner (Künzel et al., 2019) 0.519 $\pm$ 0.16
DR-learner (Kennedy, 2023b) 1.485 $\pm$ 0.18
[…]
CausalPFN (Balazadeh et al., 2025) 0.239 $\pm$ 0.11
DoPFN (Robertson et al., 2025) 0.857 $\pm$ 0.36
CausalFM (ours) 0.638 $\pm$ 0.32
Lower = better (best in bold)
H.2 Results on other datasets
In the following, we present detailed results of the experiments with ACIC 2016 datasets. We follow CausalPFN Balazadeh et al. (2025) obtaining data from https://github.com/BiomedSciAI/causallib/tree/master/causallib/datasets/data/acic_challenge_2016 to evaluate on 10 different datasets with various data generation mechanism. The treatment and outcome were simulated from real-world data corresponding to 4802 individuals and 58 covariates. Table H.1 shows the results of the CATE estimation.
H.3 Additional results for the IV setting
Table 7: IV setting for CATE estimation with binary instrument variable reported with PEHE. Results for benchmarking model performance across 10 different datasets under various confounding strength.
Method $\mathrm{D}_{1}$ $\mathrm{D}_{2}$ $\mathrm{D}_{3}$ $\mathrm{D}_{4}$ $\mathrm{D}_{5}$ $\mathrm{D}_{6}$ $\mathrm{D}_{7}$ $\mathrm{D}_{8}$ $\mathrm{D}_{9}$ $\mathrm{D}_{10}$
[…]
TARNet (Shalit et al., 2017b) 0.789 0.790 0.799 0.789 0.831 0.673 0.582 0.978 0.735 0.642
DR-learner (Kennedy, 2023b) 1.517 1.071 1.022 0.901 0.754 0.676 0.646 1.009 0.664 0.781
[…]
KIV (Singh et al., 2019) 0.660 0.344 0.340 0.394 0.544 0.460 0.299 0.731 0.532 0.241
DFIV (Xu et al., 2021) 0.654 0.245 1.022 0.459 1.145 0.770 0.741 0.366 0.971 0.717
DeepIV (Hartford et al., 2017) 0.614 0.300 0.310 0.372 0.514 0.404 0.309 0.706 0.510 0.235
DeepGMM (Bennett et al., 2019) 0.704 0.403 0.440 0.599 0.569 0.486 0.292 0.737 0.566 0.232
DMLIV (Syrgkanis et al., 2019) 0.712 0.379 0.361 0.433 0.548 0.450 0.293 0.722 0.549 0.344
DRIV (Syrgkanis et al., 2019) 0.869 0.470 0.353 0.368 0.565 0.448 0.272 0.715 0.587 0.667
MRIV (Frauen & Feuerriegel, 2022) 0.759 0.632 0.698 1.011 0.348 0.860 0.929 0.707 0.562 0.380
[…]
DoPFN (Robertson et al., 2025) 0.776 0.265 0.370 0.382 0.552 0.819 0.499 0.794 0.534 0.242
CausalFM (ours) 0.586 0.224 0.374 0.310 0.543 0.464 0.250 0.701 0.553 0.217
Reported: PEHE (mean $\pm$ standard deviation.) Lower = better (best in bold).
Table 8: IV setting for CATE estimation with continuous instrument variable reported with PEHE. Results for benchmarking model performance across 10 different datasets under various confounding strength.
Method $\mathrm{D}_{1}$ $\mathrm{D}_{2}$ $\mathrm{D}_{3}$ $\mathrm{D}_{4}$ $\mathrm{D}_{5}$ $\mathrm{D}_{6}$ $\mathrm{D}_{7}$ $\mathrm{D}_{8}$ $\mathrm{D}_{9}$ $\mathrm{D}_{10}$
[…]
TARNet (Shalit et al., 2017b) 0.943 0.825 1.025 0.458 1.007 1.316 0.848 1.004 0.825 0.884
DR-learner (Kennedy, 2023b) 1.038 1.055 0.946 0.533 0.955 1.071 1.109 1.502 1.258 0.888
[…]
KIV (Singh et al., 2019) 0.509 0.567 0.699 0.178 0.533 0.948 0.420 0.811 0.602 0.506
DFIV (Xu et al., 2021) 0.526 0.574 0.691 0.171 0.532 0.991 0.428 0.800 0.609 0.506
DeepIV (Hartford et al., 2017) 0.484 0.539 0.664 0.169 0.506 0.901 0.399 0.770 0.572 0.481
DeepGMM (Bennett et al., 2019) 0.543 0.581 0.682 0.165 0.532 1.035 0.437 0.789 0.615 0.505
DMLIV (Syrgkanis et al., 2019) 0.518 0.642 0.701 0.181 0.600 1.009 0.574 0.813 0.611 0.537
DRIV (Syrgkanis et al., 2019) 0.633 0.705 0.870 0.279 0.663 0.873 0.523 1.009 0.749 0.630
MRIV (Frauen & Feuerriegel, 2022) 0.579 0.631 0.760 0.189 0.586 1.091 0.471 0.880 0.669 0.556
[…]
DoPFN (Robertson et al., 2025) 0.471 0.528 0.787 0.322 0.649 1.723 0.416 0.588 0.722 0.545
CausalFM (ours) 0.515 0.600 0.704 0.152 0.538 0.934 0.414 0.826 0.600 0.509
Reported: PEHE (mean $\pm$ standard deviation.) Lower = better (best in bold).